Q3.A boy of mass 50 kg climbs 12steps each of height 20 cm in 50 seconds. Calculate the power used by the boy in climbing the stairs.


1
Expert's answer
2022-02-02T14:38:23-0500

By definition, the power is the work "A" divided by time "t=50s" :


"P = \dfrac{A}{t}"

The work done by the boy is:


"A = mgh"

where "m = 50kg, g=9.8m\/s^2" and "h = 12\cdot 20cm = 2.4m" is the total height. Thus, obtain:


"P = \dfrac{50kg\cdot 9.8m/s^2\cdot 2.4m}{50s} =23.52W"

Answer. 23.52W.